OpenStudy (anonymous):

Solve -2(4y + 6) = 52.

OpenStudy (anonymous):

okay, you need to distribute the -2

OpenStudy (anonymous):

Then, from there you start to solve the algebraic equation solving for why

OpenStudy (anonymous):

***y

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

hold on plz

OpenStudy (anonymous):

Okay no problem (:

OpenStudy (anonymous):

is it 8

OpenStudy (anonymous):

not quite.. not just 8

OpenStudy (anonymous):

there's something missing lol

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

Mhmm perfect (:

OpenStudy (anonymous):

\[-2(4y +6) = 52\] \[=> -8y -12 = 52\] ... \[=> -8y = 64 ..\] \[ => y= -8\]

OpenStudy (anonymous):

we open the brakets to get -4y-12=52 Like terms together -4y=52+12 so -4y=64 Devide both sides with -4 Ans y=-16